### v2

# Heavy baryons in the relativized quark model with chromodynamics

Abstract
Following the work of Capstick and Isgur [Phys. Rev. D 34, 2809 (1986)], we systematically study the mass spectrum of the heavy baryons in the relativized quark potential model with chromodynamics. Besides the original Godfrey-Isgur (GI) model, we also adopt a modified GI model which replaces the linear confinement by a screened one. The two models give similar results in our work. All heavy baryons observed so far can be explained as three-quark states. In particular, we identify the $\Omega_{c}(3000)$/$\Omega_{b}(6316)$, $\Omega_{c}(3050)$/$\Omega_{b}(6330)$, $\Omega_{c}(3065)$/$\Omega_{b}(6340)$, and $\Omega_{c}(3090)$/$\Omega_{b}(6350)$ states as the $p_{\lambda}$ excitations with quantum numbers $1/2^{-}$, $3/2^{-}$, $3/2^{-}$, and $5/2^{-}$. The $\Omega_{c}(3120)$ is a $3/2^{-}$ state with the $p_{\rho}$ excitation, whose bottom partner is predicted to be $\Omega_{b}(6446/6457,3/2^{-})$. The higher state $\Omega_{c}(3188)$ is the $2s_{\lambda}$ excitation with quantum numbers $1/2^{+}$, and $\Omega_{c}(3327)$ is a $d_{\lambda}$ excitation with quantum numbers $3/2^{+}$ or $5/2^{+}$.
I Introduction
In the past few decades, we have seen large experimental progress in searching for the heavy baryons. Lots of charmed baryons were observed in the last few years. In 2017, the LHCb collaboration observed five narrow states in the $\Xi_{c}^{+}K^{-}$ mass spectrum: $\Omega_{c}(3000)^{0}$, $\Omega_{c}(3050)^{0}$, $\Omega_{c}(3066)^{0}$, $\Omega_{c}(3090)^{0}$, and $\Omega_{c}(3119)^{0}$ [1], where the first four of them were confirmed by the Belle collaboration [2]. Later LHCb observed $\Lambda_{c}(2860)^{+}$ with quantum numbers $J^{P}=3/2^{+}$ in the $D^{0}p$ channel [3]. In 2020, LHCb further studied the $\Lambda_{c}^{+}K^{-}$ mass spectrum and observed $\Xi_{c}(2923)^{0}$, $\Xi_{c}(2939)^{0}$ and $\Xi_{c}(2965)^{0}$ [4], where the first two are new states. In 2022, the Belle collaboration observed $\Lambda_{c}(2910)^{+}$ decaying into $\Sigma_{c}\pi$ [5]. Recently, the LHCb collaboration observed two new states $\Omega_{c}(3185)^{0}$ and $\Omega_{c}(3227)^{0}$ in the $\Xi_{c}^{+}K^{-}$ decay modes [6]. In addition, many bottom baryons were also found in experiments. In 2018, the LHCb collaboration reported the observation of $\Sigma_{b}(6097)^{\pm}$ in the $\Lambda_{b}^{0}\pi^{\pm}$ final states [7]. The same year, they found the $\Xi_{b}(6227)^{-}$ as a peak in both the $\Lambda_{b}^{0}K^{-}$ and $\Xi_{b}^{0}\pi^{-}$ invariant mass spectra [8]. The isospin partner $\Xi_{b}(6227)^{0}$ was soon found [9]. In 2019, the LHCb collaboration observed the $\Lambda_{b}(6146)^{0}$ and $\Lambda_{b}(6152)^{0}$ in the $\Lambda_{b}^{0}\pi^{+}\pi^{-}$ channel [10]. Another state $\Lambda_{b}(6072)^{0}$ was also found in the same channel by the CMS collaboration[11]. In 2020, the LHCb collaboration observed four narrow peak, $\Omega_{b}(6316)^{-}$, $\Omega_{b}(6330)^{-}$, $\Omega_{b}(6340)^{-}$, and $\Omega_{b}(6350)^{-}$, in the $\Xi_{b}^{0}K^{-}$ channel [12]. Later in 2021, the CMS collaboration found $\Xi_{b}(6100)^{-}$ in the $\Xi_{b}^{-}\pi^{+}\pi^{-}$ channel [13]. Recently, the LHCb collaboration confirmed the existence of $\Xi_{b}(6227)^{0}$ and observed another state $\Xi_{b}(6333)^{0}$ [14]. To date, there are 42 singly-charmed and 25 singly-bottomed baryons collected in PDG [15], plus 3 singly-charmed and 1 singly-bottomed baryons observed more recently.
Inspired by the experimental progress, physicists used various methods to study heavy baryons, such as the quark models [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], lattice QCD [42, 43, 44, 45, 46, 47, 48], QCD sum rules [49, 50, 51, 52, 53, 54, 55], heavy quark effective theory (HQET) [56], and the Regge phenomenology [57, 58, 59]. For extra references, see recent review Refs. [60, 61, 62, 63, 64, 65].
A well-known quark potential model is the relativized quark model proposed by Godfrey and Isgur (GI) in 1985 [66], which was used to investigate the meson systems extensively [66, 67, 68, 69, 70, 71, 72]. In 1986, Capstick and Isgur extended this model to study the baryon systems [73]. Due to the lack of experimental data of heavy baryons, their work mainly focus on the light baryons and $nnQ$ baryons [Here, we assume $\mathrm{SU}(2)$ isospin symmetry and denote $n=\{u,d\}$]. In Refs. [74, 75], Lü et al. used this model, together with the diquark approximation, to study the $\Lambda_{c}$ and doubly heavy baryons. Recently,  Yu et al. studied the singly and doubly heavy baryons within this model [76, 77, 78, 79, 80]. However, they assumed $\{l_{\rho}l_{\lambda}LSj\}IJ^{P}$ ($\bm{j}=\bm{l}_{\rho}+\bm{S}$) to be good quantum numbers, while the Hamiltonian only ascertains that the $IJ^{P}$ are good quantum numbers. Furthermore, they ignored the $l_{\rho}\neq 0$ ($l_{\lambda}\neq 0$) components for the singly (doubly) heavy baryons. In this series, we will investigate the spectra of heavy baryons in the GI model, without using these approximations.
When going to the higher excited states, the coupled-channel effects may be of importance. For instance, the physical $D_{s0}^{*}(2317)$ and $D_{s1}^{*}(2460)$ are the mixture of the $c\bar{s}$ core and the $D^{(*)}K$ component [81]. Another example would be the $X(3872)$ of $J^{PC}=1^{++}$. Its mass is about 80 MeV lower than the theoretical prediction in the GI [66] and similar quark potential models. Theoretical studies suggested that the coupled-channel effect with $D\bar{D}^{*}$ is very important [82, 83, 84, 85]. In the baryon sector, the $\Lambda_{c}(2910)$ and $\Lambda_{c}(2940)$ are argued to be conventional $\Lambda_{c}$ excitations dressed with the $D^{*}N$ channel [86, 87]. In Ref. [88], Li et al. found that the coupled-channel model and the screened potential model have similar global features in describing the charmonium spectrum (below 4.0 GeV) since they approximately embody the same effect of the vacuum polarization of dynamical light quark pairs. Based on this observation, Song et al. implemented the GI model by replacing the linear confining potential $br$ by a screened one $V^{\text{scr}}=b(1-\mathrm{e}^{-{\mu}r})/\mu$ to reflect the coupled-channel effect [89]. They found that the modified GI (MGI) model is more suitable to describe the experimental data, especially the higher excited mesons [89, 90, 91, 92].
In this work, we use the two relativized quark models, both GI and MGI models, to study the mass spectra of the singly heavy baryons. In Sec II, we briefly introduce the models. Then we present the numerical results and discussions in Sec. III. Finally, we conclude in Sec. IV.
II The Relativized Quark Model
II.1 Godfrey-Isgur (GI) model
In the relativized quark model, the Hamiltonian of a baryon reads [66, 73]
$H=H_{0}+V_{\text{OGE}}+V_{\text{conf}}\,,$ (1)
where
$H_{0}=\sum_{i=1}^{3}\left(\mathbf{p}_{i}^{2}+m_{i}^{2}\right)^{1/2}$ (2)
is the relativistic kinetic energy, in which $m_{i}$ and $\mathbf{p}_{i}$ are the $i$th quark’s mass and momentum. $V_{\text{OGE}}$ is a one-gluon-exchange (OGE) potential, and $V_{\text{conf}}$ consists of a string potential and a spin-orbit term arising via Thomas precession.
Figure 1: The string confinement and the relative coordinates $\bm{\rho}$ and $\bm{\lambda}$ [73].
In the nonrelativistic limit, the potentials go to the usual Breit-Fermi interaction
$V_{\text{OGE}}\rightarrow\sum_{i<j}\left({V}_{ij}^{\text{Coul}}+{V}_{ij}^{\text{hyp}}+{V}_{ij}^{\text{so}(v)}\right)\,,$ (3)
and
$V_{\text{conf}}\rightarrow V_{\text{string}}+\sum_{i<j}{V}_{ij}^{\text{so}(v)}\,.$ (4)
Here, ${V}_{ij}^{\text{Coul}}$ is the spin-independent Coulomb-type interaction; ${V}_{ij}^{\text{hyp}}$ is the color-hyperfine interaction which consists of the contact and tensor interaction; ${V}_{ij}^{\text{so}(v)}+{V}_{ij}^{\text{so}(v)}$ is the spin-orbit interaction, and $V_{\text{string}}$ is the three-body $Y$-shape confinement interaction (see Fig. 1)
$V_{\text{string}}=C_{qqq}+b\sum_{i=1}^{3}\left|\mathbf{r}_{i}-\mathbf{r}_{\text{junction}}\right|\,.$ (5)
The $V_{\text{string}}$ interaction can be broken into an effective two-body piece and a three-body piece
$V_{\text{string}}=C_{qqq}+fb\sum_{i<j}r_{ij}+V_{3b}$ (6)
where
$V_{3b}=b\left(\sum_{i=1}^{3}\left|\mathbf{r}_{i}-\mathbf{r}_{\text{junction}}\right|-f\sum_{i<j}r_{ij}\right)$ (7)
with $f=0.5493$ [93, 94, 73] chosen to minimize the size of the expectation value of $V_{3b}$. Then the $V_{3b}$ can be treated perturbatively. Note that the contribution of $V_{3b}$ is always small [73] and we will not consider it in the present work for simplification.
As pointed out in Refs. [66, 73], the Breit-Fermi interaction should be modified for many reasons: (a) the proceeding potentials are constructed by reproducing the on-shell $qq$ scattering, which will be modified by the off-shell properties of quark; (b) the constituent quarks are not pointlike, but rather will have a graininess appropriate to some finite scale $\mu$; (c) the higher Fock space such as |$\ket{qqqg}$ is not considered in the present work, which will introduce additional momentum-dependence in the potential. To incorporate these effects, Godfrey and Isgur further built a semiquantitative model [66]. Firstly, by introducing the smearing function
$\rho_{ij}\left(\mathbf{r}-\mathbf{r}^{\prime}\right)=\frac{\sigma_{ij}^{3}}{\pi^{3/2}}\mathrm{e}^{-\sigma_{ij}^{2}\left(\mathbf{r}-\mathbf{r}^{\prime}\right)^{2}}\,,$ (8)
the potentials $V_{ij}^{\text{OGE}}$ and $V_{ij}^{\text{conf}}$ are smeared to $\tilde{V}_{ij}^{\text{OGE}}$ and $\tilde{V}_{ij}^{\text{conf}}$ via
$\tilde{V}_{ij}(r)=\int\mathrm{d}^{3}{r}^{\prime}\rho_{ij}\left(\mathbf{r}-\mathbf{r}^{\prime}\right)V_{ij}(r^{\prime}),$ (9)
with the prescription
$\sigma_{ij}^{2}=\frac{\sigma_{0}^{2}}{2}\left\{1+\left[\frac{4m_{i}m_{j}}{\left(m_{i}+m_{j}\right)^{2}}\right]^{4}\right\}+s^{2}\left(\frac{2m_{i}m_{j}}{m_{i}+m_{j}}\right)^{2},$ (10)
where $\sigma_{0}$ and $s$ are free parameters. Secondly, through the introduction of the momentum-dependent factors, the Coulomb term is modified according to
$\tilde{V}_{ij}^{\text{Coul}}{\to}\left(\beta_{ij}\right)^{1/2+\epsilon_{\text{Coul}}}\tilde{V}_{ij}^{\text{Coul}}\left(\beta_{ij}\right)^{1/2+\epsilon_{\text{Coul}}},$ (11)
and the contact, tensor, vector spin-orbit, and scalar spin-orbit potentials are modified according to
$\tilde{V}_{ij}^{\alpha}{\to}\left(\delta_{ij}^{[ij]}\right)^{1/2+\epsilon_{\alpha}}\tilde{V}_{ij}^{\alpha}\left(\delta_{ij}^{[ij]}\right)^{1/2+\epsilon_{\alpha}}$ (12)
where $\epsilon_{\alpha}$ corresponds to the contact (cont), tensor (tens), vector spin-orbit [so$(v)$], and scalar spin-orbit [$(s)$]. The momentum-dependent factors are
$\beta_{ij}=1+\frac{p_{ij}^{2}}{(p_{ij}^{2}+m_{i}^{2})^{1/2}(p_{ij}^{2}+m_{j}^{2})^{1/2}}$ (13)
and
$\delta_{ij}^{[i^{\prime}j^{\prime}]}=\frac{m_{i}m_{j}}{(p_{i^{\prime}j^{\prime}}^{2}+m_{i}^{2})^{1/2}(p_{i^{\prime}j^{\prime}}^{2}+m_{j}^{2})^{1/2}}\,,$ (14)
where $p_{ij}$ is the magnitude of the momentum of either of the quarks in the $ij$ center-of-mass frame.
Figure 2: The leading order QCD formula (15) for the effective coupling constant with $\Lambda_{\text{QCD}}=200~{}\text{MeV}$ and the fit $\alpha_{s}(Q^{2})=0.25\mathrm{e}^{-Q^{2}}+0.15\mathrm{e}^{-Q^{2}/10}+0.20\mathrm{e}^{-Q^{2}/1000}$ ($Q^{2}$ in $\text{GeV}^{2}$[66, 73].
The last piece of the model is the QCD coupling constant $\alpha_{s}$. With $N_{f}$ flavors of quark with masses below $Q^{2}$, the lowest-order coupling constant reads
$\alpha_{s}(Q^{2})=\frac{12\pi}{33-2N_{f}\ln(Q^{2}/\Lambda_{\text{QCD}}^{2})}\,.$ (15)
The $\alpha_{s}(Q^{2})$ is small for large $Q^{2}$ but diverges as $Q^{2}$ approach $\Lambda_{\text{QCD}}\approx 200~{}\text{MeV}$, which indicates the confinement. Since we are interested in the soft regime, we cannot avoid this divergence; rather we assume that $\alpha_{s}$ saturates at some value $\alpha_{s}^{\text{critical}}$ for low $Q^{2}$ as confinement emerges. For convenience, the coupling constant $\alpha_{s}$ is parametrized as
$\alpha_{s}(Q^{2})=\sum_{k}\alpha_{k}\mathrm{e}^{-Q^{2}/4\gamma_{k}^{2}}\,,$ (16)
where
$\alpha_{s}^{\text{critical}}=\sum_{k}\alpha_{k}$ (17)
is a free parameter. Fitting the QCD curve [Eq. (15), see Fig. 2] gives [66]
$\alpha_{s}(Q^{2})=0.25\mathrm{e}^{-Q^{2}}+0.15\mathrm{e}^{-Q^{2}/10}+0.20\mathrm{e}^{-Q^{2}/1000}$ (18)
with $Q^{2}$ in $\text{GeV}^{2}$.
Now we can finally present the effective potentials [73]. The smeared OGE propagator is
$\tilde{G}\left(r_{ij}\right)=-\sum_{k}\frac{2\alpha_{k}}{3r_{ij}}\cdot\text{Erf}\left(\tau_{kij}r_{ij}\right)$ (19)
where $\text{Erf}(x)$ is the Gauss error function and
$\frac{1}{\tau_{kij}^{2}}=\frac{1}{\gamma_{k}^{2}}+\frac{1}{\sigma_{ij}^{2}}\,.$ (20)
All of the OGE potential can be expressed in terms of $\tilde{G}\left(r_{ij}\right)$. More precisely, the OGE potentials
$\tilde{V}_{\text{OGE}}=\sum_{i<j}\tilde{V}_{ij}^{\text{OGE}}$ (21)
with
$\tilde{V}_{ij}^{\text{OGE}}=\tilde{V}_{ij}^{\text{Coul}}+\tilde{V}_{ij}^{\text{hyp}}+\tilde{V}_{ij}^{\text{so}(v)}$ (22)
where
$\tilde{V}_{ij}^{\text{Coul}}=\left(\beta_{ij}\right)^{1/2+\epsilon_{\text{Coul}}}\tilde{G}\left(r_{ij}\right)\left(\beta_{ij}\right)^{1/2+\epsilon_{\text{Coul}}}$ (23)
is the Coulomb-type interaction.
$\tilde{V}_{ij}^{\text{hyp}}=\tilde{V}_{ij}^{\text{cont}}+\tilde{V}_{ij}^{\text{tens}}$ (24)
is the hyperfine interaction with contact term
$\tilde{V}_{ij}^{\text{cont}}=\left(\delta_{ij}^{[ij]}\right)^{1/2+\epsilon_{\text{cont}}}\frac{2\mathbf{S}_{i}\cdot\mathbf{S}_{j}}{3m_{i}m_{j}}\nabla_{r_{ij}}^{2}\tilde{G}\left(r_{ij}\right)\left(\delta_{ij}^{[ij]}\right)^{1/2+\epsilon_{\text{cont}}}$ (25)
and tensor term
$\tilde{V}_{ij}^{\text{tens}}=\left(\delta_{ij}^{[ij]}\right)^{1/2+\epsilon_{\text{tens}}}\frac{1}{3m_{i}m_{j}}\left(\frac{3\mathbf{S}_{i}\cdot\mathbf{r}_{ij}\mathbf{S}_{j}\cdot\mathbf{r}_{ij}}{r_{ij}^{2}}-\mathbf{S}_{i}\cdot\mathbf{S}_{j}\right)\left(\frac{1}{r_{ij}}\frac{\mathrm{d}\tilde{G}\left(r_{ij}\right)}{\mathrm{d}r_{ij}}-\frac{\mathrm{d}^{2}\tilde{G}\left(r_{ij}\right)}{\mathrm{d}r_{ij}^{2}}\right)\left(\delta_{ij}^{[ij]}\right)^{1/2+\epsilon_{\text{tens}}}\,.$ (26)
$\tilde{V}_{ij}^{\text{so}(v)}={}$ $\frac{1}{r_{ij}}\frac{\mathrm{d}\tilde{G}\left(r_{ij}\right)}{\mathrm{d}r_{ij}}\Bigg{[}\left(\delta_{ii}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(v)}}\frac{\mathbf{r}_{ij}\times\mathbf{p}_{i}\cdot\mathbf{S}_{i}}{2m_{i}^{2}}\left(\delta_{ii}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(v)}}-\left(\delta_{jj}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(v)}}\frac{\mathbf{r}_{ij}\times\mathbf{p}_{j}\cdot\mathbf{S}_{j}}{2m_{j}^{2}}\left(\delta_{jj}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(v)}}$
$\qquad\qquad-\left(\delta_{ij}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(v)}}\frac{\mathbf{r}_{ij}\times\mathbf{p}_{j}\cdot\mathbf{S}_{i}-\mathbf{r}_{ij}\times\mathbf{p}_{i}\cdot\mathbf{S}_{j}}{m_{i}m_{j}}\left(\delta_{ij}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(v)}}\Bigg{]}$ (27)
is the spin-orbit interaction.
The confinement potential $V_{\text{conf}}$ consists of a string potential and a spin-orbit term arising from Thomas precession
$\tilde{V}_{\text{conf}}=C_{qqq}+\sum_{i<j}\tilde{V}_{ij}^{\text{string}}+\sum_{i<j}\tilde{V}_{ij}^{\text{so}(s)}$ (28)
with
$\tilde{V}_{ij}^{\text{string}}=fbr_{ij}\left[\frac{\mathrm{e}^{-\sigma_{ij}^{2}r_{ij}^{2}}}{\sqrt{\pi}\sigma_{ij}r_{ij}}+\left(1+\frac{1}{2\sigma_{ij}^{2}r_{ij}^{2}}\right)\text{Erf}\left(\sigma_{ij}r_{ij}\right)\right]$ (29)
and
$\tilde{V}_{ij}^{\text{so}(s)}=-\frac{1}{r_{ij}}\frac{\partial\tilde{V}_{ij}^{\text{string}}}{\partial{r}_{ij}}\Bigg{[}\left(\delta_{ii}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(s)}}\frac{\mathbf{r}_{ij}\times\mathbf{p}_{i}\cdot\mathbf{S}_{i}}{2m_{i}^{2}}\left(\delta_{ii}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(s)}}-\left(\delta_{jj}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(s)}}\frac{\mathbf{r}_{ij}\times\mathbf{p}_{j}\cdot\mathbf{S}_{j}}{2m_{j}^{2}}\left(\delta_{jj}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(s)}}\Bigg{]}$ (30)
where we have ignored the small perturbation term $V_{3b}$ [Eq. (7)].
II.2 Modified Godfrey-Isgur (MGI) model with the screening effect
In order to take into account the screening effect, Refs. [89, 90] proposed the following replacement for the linear confinement interaction
$br\rightarrow V^{\text{scr}}\left(r\right)=\frac{b\left(1-\mathrm{e}^{-\mu{r}}\right)}{\mu}\,.$ (31)
where $V^{\text{scr}}\left(r\right)$ behaves like the linear confinement potential $br$ at short distance and approaches constant $b/\mu$ at long distance. After smearing, we have
$\tilde{V}_{\text{conf}}\to\tilde{V}_{\text{conf}}^{\text{scr}}=C_{qqq}+\sum_{i<j}\tilde{V}_{ij}^{\text{scr}}+\sum_{i<j}\tilde{V}_{ij}^{\text{so(scr)}}$ (32)
where
$\tilde{V}_{ij}^{\text{scr}}={}$ $\frac{fb}{\mu}\Bigg{\{}1-\frac{1}{2\sigma_{ij}r_{ij}}\Bigg{[}\left({\frac{\mu}{2\sigma_{ij}}}+\sigma_{ij}r_{ij}\right)\exp\left(\frac{\mu^{2}}{4\sigma_{ij}^{2}}+\mu{r}_{ij}\right)\cdot\text{Erfc}\left({\frac{\mu}{2\sigma_{ij}}}+\sigma_{ij}r_{ij}\right)$
$\qquad\qquad\qquad~{}~{}-\left({\frac{\mu}{2\sigma_{ij}}}-\sigma_{ij}r_{ij}\right)\exp\left(\frac{\mu^{2}}{4\sigma_{ij}^{2}}-\mu{r}_{ij}\right)\cdot\text{Erfc}\left({\frac{\mu}{2\sigma_{ij}}}-\sigma_{ij}r_{ij}\right)\Bigg{]}\Bigg{\}}\,,$ (33)
and
$\tilde{V}_{ij}^{\text{so(scr)}}=-\frac{1}{r_{ij}}\frac{\partial\tilde{V}_{ij}^{\text{scr}}}{\partial{r}_{ij}}\Bigg{[}\left(\delta_{ii}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(s)}}\frac{\mathbf{r}_{ij}\times\mathbf{p}_{i}\cdot\mathbf{S}_{i}}{2m_{i}^{2}}\left(\delta_{ii}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(s)}}-\left(\delta_{jj}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(s)}}\frac{\mathbf{r}_{ij}\times\mathbf{p}_{j}\cdot\mathbf{S}_{j}}{2m_{j}^{2}}\left(\delta_{jj}^{[ij]}\right)^{1/2+\epsilon_{\text{so}(s)}}\Bigg{]}\,.$ (34)
where $\text{Erfc}(x)$ is the complementary error function.
II.3 Baryon wave function
To investigate the mass spectra of the baryons, we need to construct the wave functions. The total wave function is a direct product of the flavor, color, spin, and spatial wave functions.
[…]
The baryons should be color-singlet, thus we have only one totally antisymmetric color wave function
$\varphi^{c}=\ket{(q_{1}q_{2})^{3_{c}}q_{3}}^{1_{c}}\,.$ (35)
where the superscripts are color representations.
Next we consider the spin wave functions. The total spin of the baryons can be either of $1/2$ or $3/2$. A complete set of the spin wave functions are listed as follows
$\left\{\begin{split}&\chi_{1}^{s}=\ket{(q_{1}q_{2})_{1}q_{3}}_{3/2}\,,\\ &\chi_{2}^{s}=\ket{(q_{1}q_{2})_{1}q_{3}}_{1/2}\,,\\ &\chi_{3}^{s}=\ket{(q_{1}q_{2})_{0}q_{3}}_{1/2}\,,\end{split}\right.$ (36)
where the subscripts are spins $\{S_{12},S\}$. The third component $M_{S}$ is not shown for simplicity.
Finally for spatial wave functions we adopt the harmonic-oscillator wave function
$\Psi_{LM_{L}n_{\rho}l_{\rho}n_{\lambda}l_{\lambda}}\left(\bm{\rho},\bm{\lambda}\right)=\sum_{m_{l_{\rho}}m_{l_{\lambda}}}\Braket{l_{\rho}m_{l_{\rho}};l_{\lambda}m_{l_{\lambda}}}{LM_{L}}$
×$\qquad\qquad\qquad\times\psi_{n_{\rho}l_{\rho}m_{l_{\rho}}}\left(\alpha,\bm{\rho}\right)\psi_{n_{\lambda}l_{\lambda}m_{l_{\lambda}}}\left(\alpha,\bm{\lambda}\right)$ (37)
in the two relative coordinates (see Fig. 1)
$\left\{\begin{split}&\bm{\rho}=\frac{\mathbf{r}_{1}-\mathbf{r}_{2}}{\sqrt{2}}\,,\\ &\bm{\lambda}=\frac{\mathbf{r}_{1}+\mathbf{r}_{2}-2\mathbf{r}_{3}}{\sqrt{6}}\,,\end{split}\right.$ (38)
and $\psi_{nlm}(\alpha,\bm{r})$ is the harmonic oscillator wave function with variational parameter $\alpha$.
We then expand the wave function of quantum numbers $J^{P}$ in states of the form
|$\Ket{\alpha}={}$ $\chi^{f}\varphi^{c}\sum_{M_{L}M_{S}}\Braket{LM_{L};SM_{S}}{JM}$
×$\times\Psi_{LM_{L}n_{\rho}l_{\rho}n_{\lambda}l_{\lambda}}\chi_{SM_{S}}^{s}\,,$ (39)
where $P=(-1)^{l_{\rho}+l_{\lambda}}$. Note that there are infinite number of bases for any quantum numbers $IJ^{P}$. We truncate the bases by $N\leq 10$ with $N=2(n_{\rho}+n_{\lambda})+l_{\rho}+l_{\lambda}$. Diagonalizing the Hamiltonian in these bases and minimizing each eigenvalues by varying $\alpha$, we obtain the mass spectra of the heavy baryons. More details can be found in Ref. [73].
III Numerical results
III.1 Parameters
Table 1: The parameters of the relativized quark potential model. Note that we ignore isospin violation.
Parameters GI85 [66] CI86 [73] This work
GI MGI
Masses (MeV)
$m_{n}$ $220$ $220$ $247$ $244$
$m_{s}$ $419$ $419$ $455$ $452$
$m_{c}$ $1628$ $1628$ $1652$ $1644$
$m_{b}$ $4977$ $4977$ $5000$ $4991$
Potentials
$b$ ($\text{GeV}^{2}$) $0.18$ $0.15$ $0.14$ $0.17$
$\alpha_{s}^{\text{critical}}$ $0.60$ Same Same Same
$\Lambda_{\text{QCD}}$ (MeV) $220$ Same Same Same
$C_{q\bar{q}}$ (MeV) $\frac{4}{3}(-253)=-340$
$C_{qqq}$ (MeV) $-615$ $-547$ $-656$
$\mu$ (MeV) $0.06$
Relativistic effects
$\sigma_{0}$ (GeV) $1.80$ $1.80$ $1.60$ $1.60$
$s$ $1.55$ $1.55$ $1.75$ $1.67$
$\epsilon_{\text{Coul}}$ 00 $+0.060$ $+0.071$
$\epsilon_{\text{cont}}$ $-0.168$ $-0.168$ $-0.179$ $-0.173$
$\epsilon_{\text{tens}}$ $+0.025$ $-0.168$ $-0.015$ $-0.125$
$\epsilon_{\text{so}(v)}$ $-0.035$ 00 $-0.220$ $-0.237$
$\epsilon_{\text{so}(s)}$ $+0.055$ $+0.300$ $+0.039$ $-0.062$
Table 2: Masses of light and singly-heavy baryons used for fitting parameters. All masses are in units of MeV.
State, $J^{P}$ CI86 [73] This work Exp. [15] State, $J^{P}$ CI86 [73] This work Exp. [15]
GI MGI GI MGI
$N\frac{1}{2}^{+}$ 960 948 946 938.9187544±0.0000003plus-or-minus938.91875440.0000003938.9187544$\pm$ 0.0000003938.9187544 ± 0.0000003111The baryon masses are taken by their isospin averages. $\Delta\frac{3}{2}^{+}$ 1230 1233 1232 $1232\pm 2$
$\Lambda\frac{1}{2}^{+}$ 1115 1122 1122 $1115.683\pm 0.006$
$\Sigma\frac{1}{2}^{+}$ 1190 1179 1179 $1193.15\pm 0.03$ $\Sigma\frac{3}{2}^{+}$ 1370 1387 1388 $1384.6\pm 0.4$
$\Xi\frac{1}{2}^{+}$ 1305 1318 1318 $1318.28\pm 0.11$ $\Xi\frac{3}{2}^{+}$ 1505 1532 1535 $1533.4\pm 0.3$
$\Omega\frac{3}{2}^{+}$ 1635 1668 1670 $1672.45\pm 0.29$
$\Lambda_{c}\frac{1}{2}^{+}$ 2265 2290 2288 $2286.46\pm 0.14$ $\Lambda_{c}^{*}\frac{3}{2}^{+}$ 2910 2870 2871 $2856.1_{-6.0}^{+2.3}$
$\Lambda_{c}^{*}\frac{1}{2}^{-}$ 2630222Most heavy baryons are observed after the publication of Ref. [73]. These states are indicated by an underline. 2594 2592 $2592.25\pm 0.28$ $\Lambda_{c}^{*}\frac{3}{2}^{-}$ 2640 2624 2627 $2628.11\pm 0.19$
$\Sigma_{c}\frac{1}{2}^{+}$ 2440 2439 2437 $2453.45\pm 0.10$ $\Sigma_{c}\frac{3}{2}^{+}$ 2495 2519 2518 $2518.1\pm 0.3$
$\Xi_{c}\frac{1}{2}^{+}$ 2476 2474 $2469.08\pm 0.18$ $\Xi_{c}\frac{3}{2}^{+}$ 2649 2649 $2645.63\pm 0.20$
2572 2571 $2578.5\pm 0.4$
$\Xi_{c}^{*}\frac{1}{2}^{-}$ 2787 2793 $2792.9\pm 0.4$ $\Xi_{c}^{*}\frac{3}{2}^{-}$ 2814 2818 $2818.15\pm 0.20$
$\Omega_{c}\frac{1}{2}^{+}$ 2695 2692 $2695.2\pm 1.7$ $\Omega_{c}\frac{3}{2}^{+}$ 2769 2768 $2765.9\pm 2.0$
$\Lambda_{b}\frac{1}{2}^{+}$ 5585 5625 5623 $5619.60\pm 0.17$
$\Lambda_{b}^{*}\frac{1}{2}^{-}$ 5912 5894 5896 $5912.19\pm 0.17$ $\Lambda_{b}^{*}\frac{3}{2}^{-}$ 5920 5906 5908 $5920.09\pm 0.17$
$\Sigma_{b}\frac{1}{2}^{+}$ 5795 5809 5809 $5813.10\pm 0.18$ $\Sigma_{b}\frac{3}{2}^{+}$ 5805 5838 5838 $5832.53\pm 0.20$
$\Xi_{b}\frac{1}{2}^{+}$ 5805 5803 $5794.5\pm 0.4$ $\Xi_{b}\frac{3}{2}^{+}$ 5963 5964 $5953.8\pm 0.3$
5935 5934 $5935.02\pm 0.05$
$\Omega_{b}\frac{1}{2}^{+}$ 6049 6047 $6045.2\pm 1.2$
The GI model is successful in studying the mesons and baryons. In 1985, Godfrey and Isgur fitted the parameters for mesons [66]. Later Capstick and Isgur refitted the parameters for the baryons [73]. The concrete values of these parameters are collected in Table 1. After their works, lots of heavy baryons had been observed in experiments [15]. To appreciate the experimental progress in the past three decades, we perform the $\chi^{2}$ fits to obtain the 13/14 parameters of the GI/MGI models. The $\chi^{2}$ is defined as
$\chi^{2}=\sum_{i}\frac{\left(m_{i}^{\text{exp.}}-m_{i}^{\text{th.}}\right)^{2}}{\sigma_{i}^{2}}\,,$ (40)
where the $m_{i}^{\text{exp.}}$ and $m_{i}^{\text{th.}}$ are experimental and theoretical masses of the baryons. The baryons used for fits include 24 ground states and 6 excited states (see Table 2). The $\sigma_{i}$ is usually chosen to be the uncertainty of the experimental mass ${\Delta}m_{i}^{\text{exp.}}$. However, the ${\Delta}m_{N}^{\text{exp.}}\sim 10^{-7}~{}\text{MeV}$ is too small compares to the uncertainties of other baryons, which will deteriorate the fits. Thus we choose $\sigma_{i}$ to be a constant
$\sigma_{i}=k~{}\text{MeV}\,.$ (41)
The results are independent of the concrete value of $k$. Our fits give $\chi^{2}=\{1776,1558\}/k^{2}$ for the GI and MGI models, respectively. The parameters obtained are listed in Table 1. The comparison of the fitted masses with experimental data is listed in Table 2. With parameters obtained, we can calculate the heavy baryon masses. The spectroscopic results are given in Tables 36 and Figs. 310. Within uncertainty of the quark model, the GI and MGI models give similar results.
III.2 The $\Sigma_{Q}$ and $\Lambda_{Q}$ baryons
The $\Sigma_{Q}$ and $\Lambda_{Q}$ baryon masses are listed in Tables 34. Their relative positions are plotted in Figs. 36.
First we consider the $\Sigma_{c}$ baryons. From Fig. 3, we can clearly see that the low-lying states form band structures. The lowest band consists of the ground states of spin $1/2^{+}$ and $3/2^{+}$. The next band consists of seven negative parity states dominated by $P$-wave excitation components. Among them, the $1/2^{-}$ and $3/2^{-}$ states further form a two semi-band structure. The lower two states are dominated by $p_{\lambda}$ modes, while the higher ones are dominated by $p_{\rho}$ excitations. In 2004, the Belle Collaboration observed the $\Sigma_{c}(2800)$ state in the $\Lambda_{c}\pi$ channel [95]. Its quantum numbers are not determined yet [15]. Our calculation suggests that it is most likely a $3/2^{-}$ state dominated by the $p_{\lambda}$ excitation. The third band is dominated by the $N=2$ excitation (For convenience, we use the oscillator band to label the quark model states. When we refer to a state within $N=2$ band, we are implying that most part of its wave functions belong to the $N=2$ band of oscillator). More precisely, this band consists of $2s_{\rho}$, $2s_{\lambda}$, $d_{\rho}$, $d_{\lambda}$ and $[p_{\rho}p_{\lambda}]_{L}$ excitations. We again find that the excitation over ${\lambda}$ mode is easier than over ${\rho}$ mode. Taking $1/2^{+}$ states as an example, the $\Sigma_{c}(2925/2923)$ and $\Sigma_{c}(3025/3032)$ are dominated by the $2s_{\lambda}$ and $d_{\lambda}$ modes, the $\Sigma_{c}(3096/3112)$ and $\Sigma_{c}(3119/3118)$ are dominated by $d_{\rho}$ and $2s_{\rho}$ modes, while the $\Sigma_{c}(3025/3040)$ and $\Sigma_{c}(3165/3180)$ are both dominated by the $[p_{\rho}p_{\lambda}]_{L}$ mode. We also calculate the heavier states which include higher excitations. They are more packed with each other. A more detailed study of their decay properties is needed.
[…]
III.3 The $\Omega_{Q}$ baryons
Next we consider the $\Omega_{Q}$ baryons. We present their masses in Table 6 and plot their relative position in Figs. 910. The $\Omega_{Q}$ system is the strange counterpart of the $\Sigma_{Q}$ system. So their mass spectra share a similar pattern.
For the $\Omega_{c}$ system (see Fig. 9), the $\Omega_{c}$ and $\Omega_{c}^{*}$ are ground states, which are used for fits in this work. Next we consider the $P$-wave excitation with negative parity, namely the seven states lying in the $3.0\sim 3.15~{}\mathrm{GeV}$ mass region. Among them, the $\Omega_{c}(3097/3103,1/2^{-})$ and $\Omega_{c}(3132/3141,3/2^{-})$ states are dominated by the $p_{\rho}$ excitations, while the others are $p_{\lambda}$ one. As shown in Fig. 9, these states may correspond to the five narrow $\Omega_{c}$ excited states observed by the LHCb collaboration [1]. Note that the $\Omega_{c}(3020/3025,1/2^{-})$ and $\Omega_{c}(3033/3041,1/2^{-})$ states are dominated by the mixing of
$\psi_{1}=\ket{\left(1s_{\rho},1p_{\lambda}\right)_{P},1,0,1/2^{-}}\,,$ (42)
and
$\psi_{2}=\ket{\left(1s_{\rho},1p_{\lambda}\right)_{P},1,1,1/2^{-}}$ (43)
bases in the notation |$\ket{(l_{\rho},l_{\lambda})_{L},S_{12},j_{q}=L+S_{12},J^{P}}$. The calculations based on ${}^{3}P_{0}$ model suggested that a pure $\psi_{1}$ state is very broad [$\Gamma\sim\mathcal{O}(10^{2}~{}\text{MeV})$] while a pure $\psi_{2}$ state is very narrow ($\Gamma\sim 0~{}\text{MeV}$[99]. Since our parameters are solely fitted from the masses, we are unable to resolve the mixing angle between these two states with small mass difference. We can only say that one of the two states is broad. More precisely, the $\Omega_{c}(3000)$ is a $1/2^{-}$ state dominated by the $p_{\lambda}$ excitation, though we cannot determine whether it corresponds to $\Omega_{c}(3020/3025,1/2^{-})$ or $\Omega_{c}(3033/3041,1/2^{-})$. The $\Omega_{c}(3050)$ and $\Omega_{c}(3065)$ correspond to the $\Omega_{c}(3054/3059,3/2^{-})$ and $\Omega_{c}(3073/3080,3/2^{-})$, which are also dominated by the $p_{\lambda}$ excitations. The $\Omega_{c}(3090)$ is close to $\Omega_{c}(3097/3103,1/2^{-})$ and $\Omega_{c}(3104/3110,5/2^{-})$ states. However, Ref. [99] suggested that the $J^{P}=1/2^{-}$ $p_{\rho}$ excitation should be a broad state ($\sim 10^{2}~{}\text{MeV}$). Thus we interpret $\Omega_{c}(3090)$ as the $\Omega_{c}(3104/3110,5/2^{-})$ state, which is dominated by the $p_{\lambda}$ excitation . The next state, $\Omega_{c}(3120)$, corresponds to the $p_{\rho}$-dominated excitation $\Omega_{c}(3132/3143,3/2^{-})$. In the higher mass region, the LHCb collaboration found a broad structure around 3188 MeV [1], which was later confirmed by the Belle [2] and LHCb [6] collaborations. In Ref. [6], the LHCb collaboration also found the $\Omega_{c}(3327)$. Comparing their masses with our calculated masses, we interpret the $\Omega_{c}(3188)$ as the $\Omega_{c}(3176/3178,1/2^{+})$ state dominated by the $2s_{\lambda}$ excitation, while $\Omega_{c}(3327)$ is a $d_{\lambda}$-dominated excitation with quantum numbers $3/2^{+}$ or $5/2^{+}$.
Now we consider the $\Omega_{b}$ system (see Fig. 10). First we consider the ground states. Only the $1/2^{+}$ state is observed in experiment, while the $3/2^{-}$ state, $\Omega_{b}^{*}$, is not observed yet. Our calculation predicts the mass of $\Omega_{b}^{*}$ to be 6078/6077 MeV. Next we consider the $P$-wave excitations. Unlike the $\Omega_{c}$ case, an explicit two semi-band structure appears among the $P$-wave $\Omega_{b}$ excited states. The five lower states are dominated by $p_{\lambda}$ excitations while the two higher states are dominated by $p_{\rho}$ excitations. In 2020, the LHCb collaboration observed four narrow excited $\Omega_{b}$ states, namely the $\Omega_{b}(6316)$, $\Omega_{b}(6330)$, $\Omega_{b}(6340)$, and $\Omega_{b}(6350)$ states [12]. As shown in Fig. 10, the masses of these states lie in the mass region of the $p_{\lambda}$-dominated excitations. Similar to the logic of the $\Omega_{c}$ baryons, the $\Omega_{b}(6316)$ is a $1/2^{-}$ state. The $\Omega_{b}(6330)$ and $\Omega_{b}(6340)$ correspond to $\Omega_{b}(6341/6347,3/2^{-})$ and $\Omega_{b}(6355/6360,3/2^{-})$. And the $\Omega_{b}(6350)$ corresponds to $\Omega_{b}(6369/6374,5/2^{-})$. The $p_{\rho}$-dominated excitations is about 100 MeV above the $p_{\lambda}$-dominated excitations. In particular, the $\Omega_{b}(6446/6457,3/2^{-})$, which is the bottom counterpart of the $\Omega_{c}(3120)$ state, should be a narrow state. We hope the future experiments can search for these higher states.
III.4 The $\Xi_{Q}$ baryons
Now we move to the $\Xi_{Q}$ system with strangeness $-1$. Their masses are listed in Table 5. The $\Xi_{Q}$ baryons are composed of three different quarks, so they are not constrained by the Pauli principle. The consequence is that more bases are allowed to be mixed. In Figs. 78, we plot the calculated masses of the $\Xi_{Q}$ baryons. We also plot the experimental states for comparison. For the $\Xi_{c}$ baryons, we used the ground states and the lowest $1/2^{-}$ and $3/2^{-}$ states for parameter fits in Sec. III.1. There are still six excited states whose quantum numbers are undetermined. From Fig. 7, we find that it is unlikely to uniquely determine their quantum numbers. The $\Xi_{c}(2923)$/$\Xi_{c}(2930)$ might be a $1/2^{-}$ or $3/2^{-}$ state, while the $\Xi_{c}(2970)$ might be a $1/2^{-}$, $3/2^{-}$ or $5/2^{-}$ state. In the higher mass region, the quantum numbers of $\Xi_{c}(3055)$/$\Xi_{c}(3123)$ might be $1/2^{+}$ or $3/2^{+}$, while the quantum numbers of $\Xi_{c}(3030)$ might be $5/2^{\pm}$. The $\Xi_{b}$ baryons are similar. We used the three ground states for fits. For the higher states, the $\Xi_{b}(6100)$ is most likely a $3/2^{-}$ excitation, the $\Xi_{b}(6227)$ could be a $1/2^{+}$ or $3/2^{-}$ excitation, while the quantum numbers of the $\Xi_{b}(6333)$ might be $5/2^{+}$, $1/2^{-}$ or $3/2^{-}$. We hope the future experiments can determine their quantum numbers. A detailed study of their decay properties are also necessary.
IV Summary
In this work, we have systematically studied the masses of the heavy baryons in the relativized quark model, namely the (modified) Godfrey-Isgur [(M)GI] model. The key ingredients of the GI model are a universal one-gluon-exchange interaction and a linear confinement potential motivated by QCD. The relativistic effects are included semi-quantitatively. In the MGI model, the linear confinement interaction is replaced by a screened one. The parameters are fitted from the masses of the ground state baryons, the lowest $1/2^{-}$ and $3/2^{-}$ singly heavy baryons and the lowest $\frac{3}{2}^{+}\Lambda_{Q}$ states. We found that all heavy baryons observed can be explained as three-quark states. In particular,
  1. (a)
    The $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3065)$ and $\Omega_{c}(3090)$ states are $p_{\lambda}$-dominated excitations with quantum numbers $1/2^{-}$, $3/2^{-}$, $3/2^{-}$ and $5/2^{-}$, while the $\Omega_{c}(3120)$ is a $3/2^{-}$ state dominated the $p_{\rho}$ excitation. The higher state $\Omega_{c}(3188)$ is the $2s_{\lambda}$ excitation with quantum numbers $1/2^{+}$, and $\Omega_{c}(3327)$ is a $d_{\lambda}$ excitation with quantum numbers $3/2^{+}$ or $5/2^{+}$.
  2. (b)
    The $\Omega_{b}(6316)$, $\Omega_{b}(6330)$, $\Omega_{b}(6340)$, and $\Omega_{b}(6350)$ states are $p_{\lambda}$-dominated excitations with quantum numbers $1/2^{-}$, $3/2^{-}$, $3/2^{-}$, and $5/2^{-}$, while a narrow $p_{\rho}$-dominated excited state $\Omega_{b}(6446/6457,3/2^{-})$ is predicted to be about 100 MeV above them.
  3. (c)
    The $\Sigma_{c}(2800)$ is most likely a $3/2^{-}$ state dominated by the $p_{\lambda}$ excitation, while the $\Sigma_{b}(6097)$ is a $1/2^{-}$ or $3/2^{-}$ state dominated by the $p_{\lambda}$ excitation.
  4. (d)
    Our calculation reproduced the masses of the well-known $\Lambda_{c}(2880)$ and $\Lambda_{c}(2940)$ states, whose quantum numbers are $5/2^{+}$ and $3/2^{-}$ respectively. The $\Lambda_{c}(2765)$ might be a $1/2^{+}$ or $1/2^{-}$ state, and the $\Lambda_{c}(2910)$ could be a $1/2^{+}$ or $3/2^{+}$ one. Furthermore, the $\Lambda_{b}(6070)$ is a $1/2^{+}$ state dominated by the $2s_{\lambda}$ excitation, while the $\Lambda_{b}(6146)$ and $\Lambda_{b}(6152)$ states might belong to the $\{3/2^{+},5/2^{+}\}$ or $\{1/2^{-},3/2^{-}\}$ doublet.
Hopefully, our calculations will be useful for the search of the new baryon states. More experimental and theoretical investigations are expected to verify and understand the baryon system in the future.
Acknowledgments
We are grateful to Prof. Marek Karliner, Prof. Guang-Juan Wang and Dr. Lu Meng for helpful comments and discussions. We thank the High-performance Computing Platform of Peking University for providing computational resources. This project was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 11975033 and No. 12070131001; and the NSFC-ISF under Grant No. 3423/19.
Table 3: Masses of $\Lambda_{c}$ and $\Sigma_{c}$ baryons (in units of MeV).
State $J^{P}$ Model Mass
$\Lambda_{c}$ $1/2^{+}$ GI 2290 2772 2921 3007 3041 3101 3155 3185 3326 3389
MGI 2288 2772 2902 3014 3041 3135 3164 3173 3306 3379
$3/2^{+}$ GI 2870 2941 3054 3094 3169 3196 3218 3238 3338
MGI 2871 2926 3066 3117 3177 3197 3208 3237 3319
$5/2^{+}$ GI 2887 3120 3186 3228 3246 3263
MGI 2888 3140 3184 3217 3239 3273
$7/2^{+}$ GI 3219 3322
MGI 3214 3308
$9/2^{+}$ GI 3324
MGI 3311
$1/2^{-}$ GI 2594 2622 2767 3022 3115 3205 3260 3298 3323 3350 3359 3390
MGI 2592 2599 2782 3016 3098 3208 3254 3292 3324 3337 3355 3383
$3/2^{-}$ GI 2624 2826 2916 3036 3202 3235 3267 3287 3336 3349 3368
MGI 2627 2849 2916 3028 3192 3236 3266 3281 3335 3343 3367
$5/2^{-}$ GI 2973 3119 3216 3313 3350 3367
MGI 2973 3114 3209 3312 3361 3364
$7/2^{-}$ GI 3126 3355
MGI 3122 3358
$\Sigma_{c}$ $1/2^{+}$ GI 2439 2925 3025 3035 3096 3119 3165 3323 3377 3404 3457 3488
MGI 2437 2923 3032 3040 3112 3118 3180 3305 3364 3383 3431 3480
$3/2^{+}$ GI 2519 2973 3030 3044 3124 3139 3151 3176 3193 3348 3376 3389 3473 3480 3492
MGI 2518 2969 3027 3054 3135 3143 3168 3189 3217 3328 3354 3377 3453 3457 3477
$5/2^{+}$ GI 3040 3091 3165 3227 3233 3384 3418 3478 3485
MGI 3036 3092 3181 3220 3248 3360 3394 3458 3476
$7/2^{+}$ GI 3109 3280 3431 3460 3485
MGI 3109 3269 3406 3440 3479
$9/2^{+}$ GI 3459
MGI 3440
$1/2^{-}$ GI 2751 2764 2845 3158 3166 3246 3290 3304 3329 3394 3401 3423 3451
MGI 2753 2773 2852 3149 3162 3244 3287 3305 3338 3387 3410 3433 3480
$3/2^{-}$ GI 2787 2820 2890 3177 3197 3268 3273 3308 3325 3366 3370 3408 3420 3442 3447 3459 3486
MGI 2788 2825 2905 3167 3189 3263 3276 3300 3314 3371 3380 3404 3422 3452 3455 3485 3489
$5/2^{-}$ GI 2857 3223 3258 3278 3348 3372 3391 3431 3440 3456 3496
MGI 2861 3212 3248 3284 3334 3382 3395 3437 3445 3466 3487
$7/2^{-}$ GI 3261 3317 3392 3450 3463
MGI 3251 3309 3395 3436 3472
$9/2^{-}$ GI 3324 3483
MGI 3315 3464
Table 4: Masses of $\Lambda_{b}$ and $\Sigma_{b}$ baryons (in units of MeV).
State $J^{P}$ Model Mass
$\Lambda_{b}$ $1/2^{+}$ GI 5625 6047 6214 6325 6357 6406 6429 6472 6572 6659 6661
MGI 5623 6047 6195 6339 6353 6406 6457 6483 6555 6636 6660
$3/2^{+}$ GI 6130 6225 6342 6390 6450 6478 6528 6552 6579 6670 6693
MGI 6130 6207 6357 6412 6442 6488 6524 6562 6568 6670 6672
$5/2^{+}$ GI 6140 6403 6456 6493 6537 6570 6650
MGI 6141 6425 6448 6490 6534 6586 6639 6676
$7/2^{+}$ GI 6505 6525 6657
MGI 6503 6518 6647
$9/2^{+}$ GI 6530
MGI 6523
$1/2^{-}$ GI 5894 5952 6122 6271 6389 6502 6541 6578 6616 6624 6653 6679 6696
MGI 5896 5922 6141 6268 6373 6509 6528 6571 6618 6639 6648 6674 6685
$3/2^{-}$ GI 5906 6146 6274 6282 6460 6514 6545 6559 6581 6635 6641 6685
MGI 5908 6167 6273 6279 6450 6521 6532 6570 6575 6639 6653 6689
$5/2^{-}$ GI 6294 6348 6468 6570 6602 6616 6653
MGI 6295 6347 6460 6582 6588 6629 6650
$7/2^{-}$ GI 6355 6609 6620 6685
MGI 6354 6594 6633 6677
$9/2^{-}$ GI 6681 6697
MGI $6$ 6687
$\Sigma_{b}$ $1/2^{+}$ GI 5809 6237 6313 6332 6439 6446 6479 6578 6620 6661 6743 6762 6802 6837 6844 6846 6873 6880 6883
MGI 5809 6236 6324 6337 6442 6459 6502 6565 6616 6653 6726 6760 6803 6822 6834 6847 6854 6876 6883 6898
$3/2^{+}$ GI 5838 6253 6307 6322 6437 6454 6458 6491 6527 6585 6610 6627 6724 6753 6765 6777 6810 6829 6847 6856
MGI 5838 6252 6306 6334 6451 6454 6473 6519 6545 6572 6596 6622 6719 6736 6763 6771 6811 6827 6841 6852
$5/2^{+}$ GI 6314 6352 6448 6547 6582 6615 6642 6704 6730 6783 6813 6839 6840 6877 6891 6897
MGI 6313 6354 6465 6567 6573 6601 6629 6707 6723 6778 6795 6836 6845 6858 6869 6879 6887
$7/2^{+}$ GI 6362 6600 6650 6680 6709 6822 6825 6843 6897
MGI 6364 6590 6637 6669 6713 6804 6826 6849 6883 6895
$9/2^{+}$ GI 6684 6725 6828
MGI 6674 6716 6829
$1/2^{-}$ GI 6075 6081 6195 6430 6438 6535 6587 6595 6631 6696 6700 6729 6742 6750 6763 6830 6890 6898
MGI 6080 6090 6207 6424 6440 6537 6591 6597 6645 6697 6706 6723 6738 6739 6793 6818 6874 6880
$3/2^{-}$ GI 6090 6119 6212 6436 6453 6520 6544 6603 6627 6647 6659 6701 6707 6736 6746 6748 6761 6768 6788 6802
MGI 6092 6122 6227 6431 6448 6529 6545 6601 6618 6660 6668 6703 6718 6728 6740 6748 6753 6790 6797 6806
$5/2^{-}$ GI 6134 6464 6503 6527 6639 6645 6664 6711 6733 6754 6768 6784 6796 6808 6818 6846
MGI 6138 6458 6498 6536 6629 6657 6673 6728 6748 6749 6759 6763 6797 6805 6814 6850 6884 6899
$7/2^{-}$ GI 6508 6548 6651 6745 6760 6789 6825 6827 6863 6879
MGI 6504 6545 6663 6745 6760 6767 6798 6814 6868 6871 6887
$9/2^{-}$ GI 6554 6772 6830 6851 6883
MGI 6552 6757 6803 6830 6875
Table 5: Masses of $\Xi_{c}$ and $\Xi_{b}$ baryons (in units of MeV).
State $J^{P}$ Model Mass
$\Xi_{c}$ $1/2^{+}$ GI 2476 2572 2957 3054 3110 3152 3164 3166 3187 3240 3247 3252 3277 3297 3357 3447 3494 3510 3528 3531
MGI 2474 2571 2959 3055 3110 3158 3170 3176 3190 3242 3266 3278 3294 3306 3345 3434 3482 3500 3516 3522
$3/2^{+}$ GI 2649 3061 3101 3131 3173 3179 3196 3236 3257 3258 3284 3291 3299 3317 3319 3338 3410 3472 3503 3512
MGI 2649 3066 3100 3133 3175 3189 3207 3255 3261 3272 3298 3305 3313 3318 3339 3345 3398 3457 3492 3499
$5/2^{+}$ GI 3074 3181 3217 3260 3284 3300 3337 3344 3350 3370 3419 3520 3545 3587 3596
MGI 3079 3184 3222 3275 3297 3304 $3$ 3351 3364 3381 3407 3505 3529 3571 3590 3595
$7/2^{+}$ GI 3233 3330 3384 3501 3558 3596
MGI 3238 3329 3378 3494 3539 3591 3595
$9/2^{+}$ GI 3501
MGI 3494 3589
$1/2^{-}$ GI 2787 2827 2890 2901 2918 2978 3201 3286 3293 3303 3346 3372 3396 3416 3432 3435 3453 3461 3481 3503
MGI 2793 2820 2894 2910 2933 2982 3198 3281 3288 3302 3348 3369 3393 3418 3433 3439 3459 3469 3482 3500
$3/2^{-}$ GI 2814 2926 2949 2972 3016 3027 3213 3311 3325 3372 3380 3393 3400 3407 3415 3428 3447 3465 3477 3490
MGI 2818 2931 2957 2992 3026 3030 3210 3306 3322 3371 3382 3391 3401 3408 3417 3427 3442 3474 3487 3498
$5/2^{-}$ GI 2983 3082 3299 3349 3387 3401 3417 3444 3464 3485 3492 3500 3517 3545 3546 3552 3559 3571 3579 3588
MGI 2989 3084 3300 3344 3391 3399 3425 3441 3456 3494 3497 3510 3526 3545 3552 3556 3565 3576 3582 3584
$7/2^{-}$ GI 3304 3403 3442 3488 3510 3523 3559 3564 3578 3598
MGI 3305 3401 3441 3491 3512 3521 3551 3563 3585 3592
$9/2^{-}$ GI 3448 3541 3590
MGI 3446 3533 3577
$\Xi_{b}$ $1/2^{+}$ GI 5805 5935 6231 6359 6403 6440 6457 6465 6501 6550 6562 6581 6588 6589 6601 6700 6740 6751 6784 6793
MGI 5803 5934 6234 6362 6402 6453 6462 6478 6502 6563 6573 6592 6598 6605 6619 6692 6734 6750 6778 6794
$3/2^{+}$ GI 5963 6316 6377 6413 6441 6449 6480 6527 6555 6571 6586 6602 6613 6623 6638 6641 6662 6708 6741 6744
MGI 5964 6321 6379 6413 6446 6462 6494 6549 6568 6573 6602 6611 6620 6637 6640 6660 6671 6700 6734 6739
$5/2^{+}$ GI 6324 6448 6471 6538 6564 6596 6629 6648 6659 6679 6683 6746 6763 6818 6834 6837 6850 6870 6878 6898
MGI 6329 6452 6477 6560 6578 6599 6625 6650 6678 6681 6688 6739 6755 6819 6838 6841 6847 6860 6879 6884
$7/2^{+}$ GI 6480 6609 6700 6701 6770 6814 6822 6840 6886
MGI 6486 6611 6695 6700 6762 6811 6824 6847 6887
$9/2^{+}$ GI 6705 6817 6842
MGI 6705 6815 6839
$1/2^{-}$ GI 6082 6152 6204 6210 6265 6321 6445 6558 6560 6567 6639 6658 6686 6713 6717 6731 6741 6746 6754 6784
MGI 6086 6145 6212 6219 6283 6329 6446 6556 6560 6572 6647 6660 6685 6716 6722 6727 6748 6754 6763 6781
$3/2^{-}$ GI 6092 6221 6240 6286 6336 6377 6451 6566 6577 6634 6649 6650 $6$ 6688 6692 6723 6737 6741 6753 6754
MGI 6096 6226 6246 6307 6346 6380 6452 6566 6577 6637 6657 6660 6668 6689 6700 6723 6731 6734 6758 6766
$5/2^{-}$ GI 6255 6397 6524 6587 6638 6640 6656 6701 6740 6749 6758 6764 6790 6793 6821 6827 6828 6846 6857 6859
MGI 6260 6400 6529 6586 6640 6645 6668 6710 6745 6757 6769 6770 6777 6803 6827 6830 6845 6857 6864 6869
$7/2^{-}$ GI 6530 6643 6667 6746 6764 6789 6794 6833 6855 6863 6868
MGI 6535 6645 6670 6761 6775 6781 6790 6837 6853 6866 6881
$9/2^{-}$ GI 6672 6798 6861 6876
MGI 6676 6796 6853 6870
Table 6: Masses of $\Omega_{c}$ and $\Omega_{b}$ baryons (in units of MeV).
State $J^{P}$ Model Mass
$\Omega_{c}$ $1/2^{+}$ GI 2695 3176 3290 3298 3353 3390 3412 3572 3638 3650 3692 3738 3764 3772 3784 3792
MGI 2692 3178 3296 3308 3357 3405 3426 3561 3630 3640 3677 3730 3762 3769 3777 3790
$3/2^{+}$ GI 2769 3223 3304 3310 3370 3379 3417 3420 3439 3595 3642 3647 3712 3715 3735 3749 3783 3790 3798
MGI 2768 3223 3311 3322 3374 3393 3429 3430 3459 3583 3631 3640 3694 3708 3726 3742 3781 3784 3794
$5/2^{+}$ GI 3312 3340 3401 3438 3462 3649 3667 3724 3746 3757 3765
MGI 3317 3347 3413 3440 3475 3638 3657 3712 3740 3747 3753
$7/2^{+}$ GI 3354 3482 3679 3735 3746 3781
MGI 3360 3483 3667 3731 3746 3763
$9/2^{+}$ GI 3733 3764
MGI 3728 3756
$1/2^{-}$ GI 3020 3033 3097 3421 3428 3492 3533 3551 3577 3637 3647 3663 3697 3765 3771
MGI 3025 3041 3103 3419 3428 3490 3535 3553 3585 3640 3657 3673 3717 3745 3753
$3/2^{-}$ GI 3054 3073 3132 3440 3450 3511 3536 3539 3560 3605 3637 3649 3657 3670 3675 3705 3718 3780 3788
MGI 3059 3080 3143 3437 3449 3510 3536 3546 3559 3612 3646 3653 3664 3683 3687 3722 3734 3761 3769
$5/2^{-}$ GI 3104 3472 3533 3542 3571 3620 3640 3663 3670 3680 3711 3722 3735
MGI 3110 3468 3535 3551 3568 3629 3649 3671 3676 3694 3712 3728 3749 3779
$7/2^{-}$ GI 3534 3565 3624 3664 3686 3738 3752
MGI 3536 3568 3631 3662 3694 3738 3758
$9/2^{-}$ GI 3569 3689
MGI 3571 3684
$\Omega_{b}$ $1/2^{+}$ GI 6049 6475 6562 6577 6671 6707 6722 6819 6866 6901 6966 7003 7024 7059 7062 7083
MGI 6047 6479 6575 6583 6674 6721 6741 6815 6866 6896 6958 7002 7027 7054 7059 7086 7089
$3/2^{+}$ GI 6078 6492 6563 6570 6667 6679 6714 6736 6757 6827 6863 6872 6959 6975 7004 7009 7030 7048 7071 7089
MGI 6077 6495 6571 6583 6680 6683 6726 6757 6771 6822 6860 6872 6958 6967 7003 7007 7034 7054 7068 7086
$5/2^{+}$ GI 6570 6586 6675 6773 6777 6868 6879 6949 6964 7013 7024 7056 7085 7098
MGI 6577 6593 6689 6779 6788 6865 6875 6957 6963 7012 7018 7061 7091 7094
$7/2^{+}$ GI 6594 6794 6885 6937 6953 7033 7046 7087
MGI 6602 6796 6881 6940 6961 7025 7053 7094
$9/2^{+}$ GI 6941 6956 7048
MGI 6943 6956 7054
$1/2^{-}$ GI 6326 6331 6432 6678 6682 6774 6827 6833 6857 6930 6933 6960 6982 6987 6991 7069
MGI 6333 6340 6441 6680 6687 6778 6831 6838 6868 6935 6944 6967 6973 6977 7013 7060
$3/2^{-}$ GI 6341 6355 6446 6687 6695 6769 6782 6834 6841 6869 6907 6936 6937 6954 6967 6987 6992 6994 7006 7040
MGI 6347 6360 6457 6689 6697 6781 6785 6836 6842 6880 6917 6941 6952 6964 6973 6978 6980 7015 7022 7033
$5/2^{-}$ GI 6369 6705 6762 6774 6852 6869 6911 6942 6953 6959 6998 7012 7015 7033 7045 7047
MGI 6374 6706 6768 6787 6851 6881 6921 6956 6964 6970 6985 7017 7024 7029 7038 7054
$7/2^{-}$ GI 6766 6782 6874 6955 6962 7022 7038 7050 7062
MGI 6773 6787 6885 6956 6973 7025 7029 7039 7067
$9/2^{-}$ GI 6787 6967 7055
MGI 6793 6966 7044 7098
Figure 3: Masses of the $\Sigma_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 4: Masses of the $\Sigma_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 5: Masses of the $\Lambda_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15] and Ref. [5]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 6: Masses of the $\Lambda_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 7: Masses of the $\Xi_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 8: Masses of the $\Xi_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 9: Masses of the $\Omega_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 10: Masses of the $\Omega_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).